OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

This is the same question as the other one you asked except the numbers are different.

OpenStudy (anonymous):

so how i break iht down for my anser

OpenStudy (anonymous):

y-y(intial)=m(x-x(initial))

OpenStudy (anonymous):

In other words, y- (the y value you were given) = slope( which you were given)*x- (the x value you were given)

OpenStudy (anonymous):

im till confused

OpenStudy (anonymous):

y=-6x-15 right answer

OpenStudy (anonymous):

y = -6x + 20

OpenStudy (anonymous):

IS RIGHT ANWERS OKEY

OpenStudy (anonymous):

Correct answer: y= -6x-15